Mathematics 242 – Principles of Analysis Definitions Quiz 8 – April 19, 2013

Your Name:

A) (5) Let f be a bounded function on the interval [a, b]. What does it mean for f to be integrable on [a, b]?

B) (5) We say a function f has a local maximum at c if f(c) ≥ f(x) for all x in some interval containing c. True/False and explain your answer: If f has a local maximum at c and f′(c) exists, then f′(c) = 0.